# Conduct Parameter Estimation in Homogeneous Goods Markets with Equilibrium Existence and Uniqueness Conditions: The Case of Log-linear Specification

### v1

@@ A.4 Additional experiments
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.347 2.978 -0.153 2.547 0.065 0.976 -0.009 0.760
$\alpha_{1}$0.043 0.452 -0.016 0.388 0.010 0.150 -0.001 0.117
$\alpha_{2}$0.024 0.065 -0.017 0.054 0.000 0.012 0.000 0.010
$\alpha_{3}$0.028 0.265 -0.013 0.224 0.005 0.091 0.001 0.073
$\gamma_{0}$2.458 5.660 -1.258 2.959 -0.132 0.942 -0.153 0.748
$\gamma_{1}$ 0.112 0.602 0.057 0.260 0.014 0.103 0.014 0.082
$\gamma_{2}$ 0.053 0.314 0.030 0.172 0.007 0.073 0.010 0.060
$\gamma_{3}$ 0.061 0.426 0.028 0.181 0.006 0.077 0.008 0.058
$\theta$ 0.169 0.407 0.109 0.356 -0.003 0.155 0.017 0.116
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.614 5.995 -0.213 4.315 0.077 2.034 0.063 1.555
$\alpha_{1}$0.085 0.902 -0.024 0.663 0.011 0.312 0.010 0.240
$\alpha_{2}$0.028 0.105 -0.022 0.073 0.000 0.025 0.001 0.020
$\alpha_{3}$0.070 0.549 -0.019 0.431 -0.001 0.185 0.004 0.152
$\gamma_{0}$5.106 15.922 -2.379 6.990 -0.375 1.959 -0.398 1.533
$\gamma_{1}$ 0.386 2.047 0.141 0.839 0.045 0.229 0.044 0.175
$\gamma_{2}$ 0.190 1.155 0.054 0.475 0.022 0.150 0.027 0.120
$\gamma_{3}$ 0.163 1.006 0.065 0.482 0.013 0.149 0.023 0.121
$\theta$ 0.186 0.442 0.158 0.422 -0.007 0.275 0.014 0.217
Runs converged (%) 100.000 100.000 100.000 100.000
Sample size ($T$) 100 200 1000 1500
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.835 9.908 -0.561 6.310 0.147 4.424 0.301 3.686
$\alpha_{1}$0.109 1.563 -0.076 0.935 0.024 0.677 0.047 0.566
$\alpha_{2}$0.043 0.124 -0.033 0.135 -0.005 0.061 0.001 0.045
$\alpha_{3}$0.047 0.909 -0.078 0.664 -0.004 0.404 0.018 0.323
$\gamma_{0}$6.629 20.418 -5.373 16.692 -1.445 4.738 -1.123 3.411
$\gamma_{1}$ 0.422 2.630 0.384 2.103 0.137 0.580 0.115 0.409
$\gamma_{2}$ 0.184 1.788 0.185 1.318 0.073 0.357 0.048 0.261
$\gamma_{3}$ 0.199 1.885 0.215 1.496 0.068 0.376 0.055 0.276
$\theta$ 0.225 0.459 0.227 0.447 0.098 0.357 0.084 0.313
Runs converged (%) 99.200 99.800 99.800 100.000
Sample size ($T$) 100 200 1000 1500
@@
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.018 0.465 0.007 0.323 -0.008 0.213 -0.006 0.097
$\alpha_{1}$0.045 2.257 0.024 1.523 0.018 1.016 -0.031 0.455
$\alpha_{2}$0.001 0.255 -0.001 0.176 -0.004 0.115 0.001 0.051
$\alpha_{3}$0.005 0.108 0.003 0.075 -0.001 0.050 -0.001 0.022
$\gamma_{0}$0.061 0.732 -0.005 0.474 -0.021 0.346 -0.005 0.152
$\gamma_{1}$0.311 3.450 -0.124 1.928 -0.081 1.303 -0.003 0.548
$\gamma_{2}$ 0.009 0.109 -0.001 0.071 0.003 0.051 0.000 0.023
$\gamma_{3}$ 0.001 0.108 0.003 0.075 0.003 0.053 0.000 0.022
$\theta$ 0.047 0.354 0.017 0.209 0.014 0.135 0.003 0.058
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.027 1.023 -0.002 0.641 -0.004 0.448 -0.016 0.188
$\alpha_{1}$0.024 4.396 -0.169 2.965 0.061 2.060 0.011 0.905
$\alpha_{2}$0.006 0.494 0.016 0.325 -0.007 0.234 -0.006 0.100
$\alpha_{3}$0.006 0.223 0.002 0.153 -0.001 0.099 -0.003 0.045
$\gamma_{0}$0.318 1.769 -0.091 1.059 -0.086 0.714 -0.004 0.308
$\gamma_{1}$ 5.859 210.853 -0.679 6.280 -0.338 2.972 -0.050 1.110
$\gamma_{2}$ 0.035 0.247 0.011 0.157 0.011 0.104 0.000 0.045
$\gamma_{3}$ 0.045 0.250 0.012 0.150 0.010 0.104 0.002 0.045
$\theta$0.399 18.450 0.098 0.738 0.054 0.308 0.009 0.113
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.263 2.596 0.071 1.670 -0.040 0.947 -0.002 0.412
$\alpha_{1}$0.271 10.820 0.008 6.492 0.236 4.263 0.021 1.809
$\alpha_{2}$0.044 1.253 0.023 0.779 -0.031 0.483 -0.003 0.210
$\alpha_{3}$0.024 0.584 0.008 0.343 -0.004 0.225 0.003 0.092
$\gamma_{0}$2.074 19.624 -0.551 3.043 -0.171 1.516 -0.051 0.633
$\gamma_{1}$ 58.209 1750.688 -2.416 56.909 -3.617 39.044 -0.103 2.334
$\gamma_{2}$ 0.242 2.430 0.065 0.409 0.020 0.220 0.006 0.093
$\gamma_{3}$ 0.230 2.328 0.055 0.404 0.010 0.219 0.008 0.092
$\theta$6.668 233.851 0.372 6.334 0.418 3.820 0.024 0.245
Sample size ($T$) 50 100 200 1000
Note: The data generating process follows Matsumura and Otani (2023b).
@@
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.013 0.462 0.008 0.322 -0.008 0.213 -0.006 0.097
$\alpha_{1}$0.096 2.201 0.015 1.511 0.018 1.016 -0.031 0.455
$\alpha_{2}$ 0.006 0.247 0.001 0.174 -0.004 0.115 0.001 0.051
$\alpha_{3}$0.004 0.108 0.003 0.074 -0.001 0.050 -0.001 0.022
$\gamma_{0}$0.054 0.724 -0.002 0.472 -0.021 0.346 -0.005 0.152
$\gamma_{1}$0.098 2.620 -0.093 1.847 -0.081 1.303 -0.003 0.548
$\gamma_{2}$ 0.008 0.108 -0.002 0.070 0.003 0.051 0.000 0.023
$\gamma_{3}$ 0.001 0.107 0.003 0.075 0.003 0.053 0.000 0.022
$\theta$ 0.023 0.258 0.014 0.197 0.014 0.135 0.003 0.058
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$ 0.012 1.015 0.012 0.636 0.001 0.446 -0.016 0.188
$\alpha_{1}$0.383 4.216 -0.291 2.872 0.015 2.012 0.011 0.905
$\alpha_{2}$ 0.042 0.466 0.033 0.310 -0.001 0.227 -0.006 0.100
$\alpha_{3}$ 0.000 0.222 0.004 0.152 0.000 0.099 -0.003 0.045
$\gamma_{0}$0.233 1.612 -0.067 1.020 -0.076 0.703 -0.004 0.308
$\gamma_{1}$ 0.406 4.332 -0.024 3.174 -0.178 2.585 -0.050 1.110
$\gamma_{2}$ 0.025 0.231 0.008 0.154 0.010 0.103 0.000 0.045
$\gamma_{3}$ 0.034 0.233 0.010 0.146 0.009 0.103 0.002 0.045
$\theta$ 0.014 0.380 0.018 0.311 0.035 0.260 0.009 0.113
Sample size ($T$) 50 100 200 1000
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\alpha_{0}$0.138 2.592 0.142 1.671 0.004 0.936 0.000 0.410
$\alpha_{1}$0.986 11.090 -0.665 6.325 -0.174 4.042 0.003 1.788
$\alpha_{2}$ 0.065 1.255 0.112 0.756 0.023 0.447 0.000 0.207
$\alpha_{3}$0.006 0.589 0.019 0.345 0.003 0.224 0.003 0.092
$\gamma_{0}$0.123 3.350 -0.298 2.455 -0.081 1.379 -0.047 0.628
$\gamma_{1}$ 0.635 7.424 0.141 5.668 -0.081 4.255 -0.040 2.159
$\gamma_{2}$ 0.016 0.484 0.037 0.349 0.009 0.205 0.005 0.093
$\gamma_{3}$ 0.009 0.517 0.026 0.336 0.000 0.208 0.007 0.092
$\theta$0.029 0.446 0.025 0.414 0.034 0.381 0.016 0.220
Sample size ($T$) 50 100 200 1000
Note: The data generating process follows Matsumura and Otani (2023b).

### v2

@@ A.5 Additional experiments
(a) $\sigma=0.5$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.347 2.978 -0.153 2.547 0.065 0.976 -0.009 0.760 $\alpha_{1}$ -0.043 0.452 -0.016 0.388 0.010 0.150 -0.001 0.117 $\alpha_{2}$ -0.024 0.065 -0.017 0.054 0.000 0.012 0.000 0.010 $\alpha_{3}$ -0.028 0.265 -0.013 0.224 0.005 0.091 0.001 0.073 $\gamma_{0}$ -2.458 5.660 -1.258 2.959 -0.132 0.942 -0.153 0.748 $\gamma_{1}$ 0.112 0.602 0.057 0.260 0.014 0.103 0.014 0.082 $\gamma_{2}$ 0.053 0.314 0.030 0.172 0.007 0.073 0.010 0.060 $\gamma_{3}$ 0.061 0.426 0.028 0.181 0.006 0.077 0.008 0.058 $\theta$ 0.169 0.407 0.109 0.356 -0.003 0.155 0.017 0.116 Runs converged (%) 100.000 100.000 100.000 100.000 Sample size ($T$) 100 200 1000 1500
(c) $\sigma=2.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.835 9.908 -0.561 6.310 0.147 4.424 0.301 3.686 $\alpha_{1}$ -0.109 1.563 -0.076 0.935 0.024 0.677 0.047 0.566 $\alpha_{2}$ -0.043 0.124 -0.033 0.135 -0.005 0.061 0.001 0.045 $\alpha_{3}$ -0.047 0.909 -0.078 0.664 -0.004 0.404 0.018 0.323 $\gamma_{0}$ -6.629 20.418 -5.373 16.692 -1.445 4.738 -1.123 3.411 $\gamma_{1}$ 0.422 2.630 0.384 2.103 0.137 0.580 0.115 0.409 $\gamma_{2}$ 0.184 1.788 0.185 1.318 0.073 0.357 0.048 0.261 $\gamma_{3}$ 0.199 1.885 0.215 1.496 0.068 0.376 0.055 0.276 $\theta$ 0.225 0.459 0.227 0.447 0.098 0.357 0.084 0.313 Runs converged (%) 99.200 99.800 99.800 100.000 Sample size ($T$) 100 200 1000 1500
@@
(a) $\sigma=0.5$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.018 0.465 0.007 0.323 -0.008 0.213 -0.006 0.097 $\alpha_{1}$ -0.045 2.257 0.024 1.523 0.018 1.016 -0.031 0.455 $\alpha_{2}$ -0.001 0.255 -0.001 0.176 -0.004 0.115 0.001 0.051 $\alpha_{3}$ -0.005 0.108 0.003 0.075 -0.001 0.050 -0.001 0.022 $\gamma_{0}$ -0.061 0.732 -0.005 0.474 -0.021 0.346 -0.005 0.152 $\gamma_{1}$ -0.311 3.450 -0.124 1.928 -0.081 1.303 -0.003 0.548 $\gamma_{2}$ 0.009 0.109 -0.001 0.071 0.003 0.051 0.000 0.023 $\gamma_{3}$ 0.001 0.108 0.003 0.075 0.003 0.053 0.000 0.022 $\theta$ 0.047 0.354 0.017 0.209 0.014 0.135 0.003 0.058 Sample size ($T$) 50 100 200 1000
(b) $\sigma=1.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.027 1.023 -0.002 0.641 -0.004 0.448 -0.016 0.188 $\alpha_{1}$ -0.024 4.396 -0.169 2.965 0.061 2.060 0.011 0.905 $\alpha_{2}$ -0.006 0.494 0.016 0.325 -0.007 0.234 -0.006 0.100 $\alpha_{3}$ -0.006 0.223 0.002 0.153 -0.001 0.099 -0.003 0.045 $\gamma_{0}$ -0.318 1.769 -0.091 1.059 -0.086 0.714 -0.004 0.308 $\gamma_{1}$ 5.859 210.853 -0.679 6.280 -0.338 2.972 -0.050 1.110 $\gamma_{2}$ 0.035 0.247 0.011 0.157 0.011 0.104 0.000 0.045 $\gamma_{3}$ 0.045 0.250 0.012 0.150 0.010 0.104 0.002 0.045 $\theta$ -0.399 18.450 0.098 0.738 0.054 0.308 0.009 0.113 Sample size ($T$) 50 100 200 1000
(c) $\sigma=2.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.263 2.596 0.071 1.670 -0.040 0.947 -0.002 0.412 $\alpha_{1}$ -0.271 10.820 0.008 6.492 0.236 4.263 0.021 1.809 $\alpha_{2}$ -0.044 1.253 0.023 0.779 -0.031 0.483 -0.003 0.210 $\alpha_{3}$ -0.024 0.584 0.008 0.343 -0.004 0.225 0.003 0.092 $\gamma_{0}$ -2.074 19.624 -0.551 3.043 -0.171 1.516 -0.051 0.633 $\gamma_{1}$ 58.209 1750.688 -2.416 56.909 -3.617 39.044 -0.103 2.334 $\gamma_{2}$ 0.242 2.430 0.065 0.409 0.020 0.220 0.006 0.093 $\gamma_{3}$ 0.230 2.328 0.055 0.404 0.010 0.219 0.008 0.092 $\theta$ -6.668 233.851 0.372 6.334 0.418 3.820 0.024 0.245 Sample size ($T$) 50 100 200 1000
Note: The data generating process follows matsumura2023resolving.
@@
(a) $\sigma=0.5$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.013 0.462 0.008 0.322 -0.008 0.213 -0.006 0.097 $\alpha_{1}$ -0.096 2.201 0.015 1.511 0.018 1.016 -0.031 0.455 $\alpha_{2}$ 0.006 0.247 0.001 0.174 -0.004 0.115 0.001 0.051 $\alpha_{3}$ -0.004 0.108 0.003 0.074 -0.001 0.050 -0.001 0.022 $\gamma_{0}$ -0.054 0.724 -0.002 0.472 -0.021 0.346 -0.005 0.152 $\gamma_{1}$ -0.098 2.620 -0.093 1.847 -0.081 1.303 -0.003 0.548 $\gamma_{2}$ 0.008 0.108 -0.002 0.070 0.003 0.051 0.000 0.023 $\gamma_{3}$ 0.001 0.107 0.003 0.075 0.003 0.053 0.000 0.022 $\theta$ 0.023 0.258 0.014 0.197 0.014 0.135 0.003 0.058 Sample size ($T$) 50 100 200 1000
(b) $\sigma=1.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ 0.012 1.015 0.012 0.636 0.001 0.446 -0.016 0.188 $\alpha_{1}$ -0.383 4.216 -0.291 2.872 0.015 2.012 0.011 0.905 $\alpha_{2}$ 0.042 0.466 0.033 0.310 -0.001 0.227 -0.006 0.100 $\alpha_{3}$ 0.000 0.222 0.004 0.152 0.000 0.099 -0.003 0.045 $\gamma_{0}$ -0.233 1.612 -0.067 1.020 -0.076 0.703 -0.004 0.308 $\gamma_{1}$ 0.406 4.332 -0.024 3.174 -0.178 2.585 -0.050 1.110 $\gamma_{2}$ 0.025 0.231 0.008 0.154 0.010 0.103 0.000 0.045 $\gamma_{3}$ 0.034 0.233 0.010 0.146 0.009 0.103 0.002 0.045 $\theta$ 0.014 0.380 0.018 0.311 0.035 0.260 0.009 0.113 Sample size ($T$) 50 100 200 1000
(c) $\sigma=2.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.138 2.592 0.142 1.671 0.004 0.936 0.000 0.410 $\alpha_{1}$ -0.986 11.090 -0.665 6.325 -0.174 4.042 0.003 1.788 $\alpha_{2}$ 0.065 1.255 0.112 0.756 0.023 0.447 0.000 0.207 $\alpha_{3}$ -0.006 0.589 0.019 0.345 0.003 0.224 0.003 0.092 $\gamma_{0}$ -0.123 3.350 -0.298 2.455 -0.081 1.379 -0.047 0.628 $\gamma_{1}$ 0.635 7.424 0.141 5.668 -0.081 4.255 -0.040 2.159 $\gamma_{2}$ 0.016 0.484 0.037 0.349 0.009 0.205 0.005 0.093 $\gamma_{3}$ 0.009 0.517 0.026 0.336 0.000 0.208 0.007 0.092 $\theta$ -0.029 0.446 0.025 0.414 0.034 0.381 0.016 0.220 Sample size ($T$) 50 100 200 1000
Note: The data generating process follows matsumura2023resolving.
Table 9: N2SLS with Constraints (13) and (14) for the loglinear model
(a) $\sigma=0.5$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ 0.066 3.646 0.536 3.302 0.638 2.013 0.544 1.606 $\alpha_{1}$ 0.011 0.559 0.084 0.508 0.122 0.451 0.103 0.383 $\alpha_{2}$ 0.001 0.048 0.004 0.040 0.006 0.037 0.006 0.042 $\alpha_{3}$ 0.006 0.317 0.039 0.280 0.049 0.161 0.044 0.131 $\gamma_{0}$ 8.768 9.771 10.695 11.255 12.797 12.854 12.890 12.928 $\gamma_{1}$ -0.143 0.364 -0.173 0.261 -0.182 0.203 -0.180 0.192 $\gamma_{2}$ -0.069 0.247 -0.081 0.164 -0.089 0.111 -0.086 0.099 $\gamma_{3}$ -0.067 0.239 -0.085 0.170 -0.089 0.111 -0.087 0.100 $\theta$ -19426.126 60643.314 -34723.006 57457.978 -64958.969 153375.492 -61625.065 71909.012 Runs converged (%) 100.000 100.000 96.500 93.400 Sample size ($T$) 100 200 1000 1500
(b) $\sigma=1.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.036 7.004 0.110 5.112 0.576 3.164 0.582 2.252 $\alpha_{1}$ -0.158 1.057 0.022 0.782 0.116 0.682 0.122 0.638 $\alpha_{2}$ -0.009 0.103 0.000 0.102 0.007 0.050 0.008 0.056 $\alpha_{3}$ -0.098 0.615 0.007 0.481 0.036 0.258 0.043 0.195 $\gamma_{0}$ 8.521 15.430 10.972 11.923 12.729 12.840 12.831 12.893 $\gamma_{1}$ -0.039 1.684 -0.184 0.517 -0.184 0.252 -0.176 0.222 $\gamma_{2}$ -0.016 0.977 -0.100 0.348 -0.090 0.154 -0.079 0.127 $\gamma_{3}$ -0.037 0.781 -0.096 0.303 -0.098 0.158 -0.084 0.133 $\theta$ -159519.119 1344515.437 -106885.344 627588.561 -69229.439 166381.207 -62599.755 74618.949 Runs converged (%) 99.800 99.500 97.400 93.000 Sample size ($T$) 100 200 1000 1500
(c) $\sigma=2.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -2.128 11.259 -1.140 8.691 0.744 5.721 0.778 4.198 $\alpha_{1}$ -0.325 1.767 -0.175 1.309 0.147 1.100 0.142 0.709 $\alpha_{2}$ -0.019 0.139 -0.007 0.145 0.011 0.076 0.010 0.066 $\alpha_{3}$ -0.161 0.971 -0.112 0.866 0.040 0.470 0.057 0.358 $\gamma_{0}$ 9.831 18.315 10.885 13.580 12.904 13.202 12.878 13.060 $\gamma_{1}$ -0.065 2.134 -0.125 1.066 -0.207 0.404 -0.179 0.330 $\gamma_{2}$ -0.100 1.292 -0.082 0.723 -0.090 0.275 -0.095 0.219 $\gamma_{3}$ -0.050 1.475 -0.051 0.759 -0.099 0.281 -0.088 0.229 $\theta$ -696456.922 5688252.310 -418083.860 2991591.323 -84373.063 140028.575 -80810.876 148172.098 Runs converged (%) 99.100 99.100 95.500 90.600 Sample size ($T$) 100 200 1000 1500
Table 10: N2SLS with Constraints (12) for the loglinear model
(a) $\sigma=0.5$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -0.087 3.688 0.145 3.117 0.027 0.975 -0.066 0.764 $\alpha_{1}$ -0.014 0.565 0.021 0.477 0.004 0.150 -0.010 0.117 $\alpha_{2}$ 0.001 0.045 0.003 0.036 0.000 0.012 -0.001 0.010 $\alpha_{3}$ -0.004 0.317 0.009 0.263 0.002 0.090 -0.004 0.073 $\gamma_{0}$ -0.679 3.764 -0.257 2.192 0.015 0.984 0.040 0.815 $\gamma_{1}$ 0.097 0.465 0.041 0.253 0.004 0.104 0.000 0.084 $\gamma_{2}$ 0.049 0.290 0.023 0.170 0.002 0.073 0.003 0.060 $\gamma_{3}$ 0.050 0.288 0.021 0.178 0.001 0.077 0.001 0.058 $\theta$ -0.063 0.389 -0.068 0.333 -0.043 0.199 -0.032 0.167 Runs converged (%) 99.500 99.700 100.000 100.000 Sample size ($T$) 100 200 1000 1500
(b) $\sigma=1.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -1.922 8.603 -0.068 5.116 0.037 2.035 0.000 1.556 $\alpha_{1}$ -0.299 1.314 -0.010 0.785 0.005 0.312 0.000 0.240 $\alpha_{2}$ -0.013 0.104 -0.002 0.063 0.001 0.024 0.000 0.019 $\alpha_{3}$ -0.165 0.774 -0.007 0.472 -0.004 0.185 -0.001 0.152 $\gamma_{0}$ -1.767 14.394 -1.001 6.530 -0.208 1.993 -0.156 1.566 $\gamma_{1}$ 0.255 1.949 0.132 0.838 0.034 0.229 0.027 0.174 $\gamma_{2}$ 0.125 1.097 0.053 0.475 0.017 0.150 0.019 0.119 $\gamma_{3}$ 0.099 0.903 0.062 0.481 0.007 0.149 0.014 0.120 $\theta$ -0.098 0.441 -0.060 0.421 -0.061 0.319 -0.058 0.281 Runs converged (%) 98.100 98.700 100.000 100.000 Sample size ($T$) 100 200 1000 1500
(c) $\sigma=2.0$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE $\alpha_{0}$ -4.309 16.220 -2.888 12.102 -0.143 5.650 0.085 3.693 $\alpha_{1}$ -0.657 2.489 -0.448 1.840 -0.024 0.874 0.012 0.568 $\alpha_{2}$ -0.035 0.178 -0.023 0.170 0.001 0.062 0.002 0.042 $\alpha_{3}$ -0.280 1.174 -0.250 1.181 -0.027 0.545 0.003 0.323 $\gamma_{0}$ 2.206 15.660 -0.576 13.314 -0.897 4.882 -0.646 3.317 $\gamma_{1}$ -0.255 2.096 0.100 1.753 0.119 0.607 0.086 0.399 $\gamma_{2}$ -0.163 1.316 0.045 1.113 0.065 0.363 0.034 0.257 $\gamma_{3}$ -0.155 1.669 0.068 1.127 0.061 0.386 0.041 0.271 $\theta$ -0.201 0.474 -0.142 0.463 -0.053 0.400 -0.050 0.374 Runs converged (%) 98.100 97.200 99.200 99.400 Sample size ($T$) 100 200 1000 1500